# Efficient Conditional Gradient Methods for Solving Stochastic Convex Bilevel Optimization Problems

### v1

@@ A.2 Proof of Theorem 1
Proof of Theorem 1.
This is an immediate consequence of Lemmas 5 and 6. ∎
A.3 Proof of Theorem 2
@@
The proof proceeds by applying Proposition 1 with $k=q$. Unlike the proof of Theorem 3, here we have two main random components in Eq. 11 and Eq. 12: $(q+1)q\left(\Phi_{q}(x_{q})-\Phi_{q}(x_{\operatorname{opt}})\right)$ and $\sum_{i\in[t]\setminus[q]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$. We first derive probabilistic upper bounds for these terms, then carefully set the parameters in Algorithm 2 to establish the convergence rate for both lower- and upper-level problems. In the following we use the notation
$F(x)=\sum_{i\in[n]}f_{i}(x),\quad G(x)=\sum_{i\in[n]}g_{i}(x).$
@@
We begin with a probabilistic bound on the gradient estimator used in Algorithm 2.
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\log(q)/q$ for any $0\leq t<q$ and $\alpha_{t}=2/(t+2)$ for any $t\geq q$. If 1 holds, then for any $t\geq 0$, given $\delta\in(0,1)$, with probability at least $1-\delta$, it (jointly) holds that
$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|$ $\leq 8L_{f}D\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(\frac{8}{\delta}\right)},$ (17)
$\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|$ $\leq 8L_{g}D\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(\frac{8}{\delta}\right)},$
@@
$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|$ $\leq 8L_{f}D\sqrt{\frac{(t-s_{t})}{S(t+1)(s_{t}+1)}}\sqrt{\log\left(\frac{8}{\delta}\right)}$ (18)
$\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|$ $\leq 8L_{g}D\sqrt{\frac{(t-s_{t})}{S(t+1)(s_{t}+1)}}\sqrt{\log\left(\frac{8}{\delta}\right)},$
@@
if $t>q$. Given $\delta\in(0,1)$ and $1\leq t\leq q$, setting the right hand side $=\delta/2$ and solving for $\lambda$, we have
$P\left(\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|\geq 8L_{f}D\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(\frac{8}{\delta}\right)}\right)\leq\frac{\delta}{2}$
@@
$P\left(\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|\geq 8L_{g}D\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(\frac{8}{\delta}\right)}\right)\leq\frac{\delta}{2}.$
@@
To establish convergence results for Algorithm 2, we replace 3 with the following.
@@
Note that $\sigma_{t}$ is fixed during the initialization phase, and $\sigma_{t}^{2}$ in 3 is replaced by $\sigma_{t}^{2}$ in 4. Unlike the proof of Theorem 1, where we set $k=0$, the proof of Theorem 2 requires setting $k=q$. We therefore begin by analyzing the sequence $\{x_{t}\}_{t\geq 0}$ generated by Algorithm 2 over the first $q$ iterations.
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\log(q)/q$ for any $0\leq t<q$ and $S=q$. If 1 and 4 hold, we have then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{0\leq i<q}1/(i+1)(i+2)$, for any $t\geq q$, it holds that
@@
Proof of Lemma 8.
By Lemma 1, for any $0\leq i<q$, we have
[…]
By 4, we have $\alpha_{0}=\dots=\alpha_{q-1}$ and
@@
where both inequalities are direct consequences of 4 and (19). Note that
@@
[…]
@@
Hence, applying Lemma 7 along with a union bound argument implies that, for any $\delta\in(0,1)$, with probability at least $1-\delta\sum_{0\leq i<q}1/(i+1)(i+2)$, we have
@@
Plugging the above probabilistic bound into Eq. 20, noting that $t\geq q$ and $S=q$, and using the definition of $\alpha_{0}$ complete the proof. ∎
We next analyze the sequence $\{x_{t}\}_{t\geq 0}$ generated by Algorithm 2 after the first $q$ iterations.
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\log(q)/q$ for any $0\leq t<q$, $\alpha_{t}=2/(t+2)$ for any $t\geq q$ and $S=q$. If 1 holds and the parameters $\{\sigma_{t}\}_{t\geq 0}$ are non-increasing and positive, then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{i\in[t]\setminus[q]}1/i(i+1)$, for any $t>q$, it holds that
@@
Given $\delta\in(0,1)$ and $q<i\leq 2q$, by Lemma 7 and the inequality Eq. 22, with probability at least $1-\delta/i(i+1)$, we have
@@
where the second inequality holds as $\frac{1}{q}\sum_{q<i\leq t}i\leq\sqrt{2}t-q/2$ for all $q<t\leq 2q$. Similarly, given $\delta\in(0,1)$ and $i>2q$, by Lemma 7 and the inequality Eq. 22, with probability at least $1-\delta/i(i+1)$, we have
@@
where the second inequality follows since $q\lfloor(i-1)/q\rfloor+1\geq i-q$, and the third inequality follows since $i/(i-q)<2$ when $i>2q$. Thus, if $t>2q$, with probability at least $1-\delta\sum_{2q<i\leq t}1/i(i+1)=1-\delta$, we have
@@
Lemma 16 in Section B.4 guarantees that these quantities remain finite for any parameter choice satisfying 1 to 2 and 4. In particular, these conditions hold when $\sigma_{t}=\varsigma(t+1)^{p}$ for any $p\in(0,1)$, which is what is used in Theorem 2. We now analyze the sequence $\{x_{t}\}_{t\geq 0}$ in Algorithm 2 for the inner-level problem.
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\log(q)/q$ for $0\leq t<q$ $\alpha_{t}=2/(t+2)$ for any $t\geq q$, $S=q$, and let $C_{\delta,q}$ be defined as in (23g). If 1 and 1, 2 and 4 hold, then with probability at least $1-\delta$, it holds that
@@
Proof of Lemma 9.
Combining (11) from Proposition 1 with the probabilistic bounds in Lemma 8 and Proposition 3, and after some straightforward calculations, with probability at least $1-\delta$, we have
@@
We next analyze the convergence of the sequence $\{z_{t}\}_{t\geq 0}$ in Algorithm 2 for both outer- and inner-level problems.
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 2 with $\alpha_{t}=\log(q)/q$ for $0\leq t<q$ $\alpha_{t}=2/(t+2)$ for any $t\geq q$, $S=q$, and let $C_{\delta,q}$ and $V_{q}$ defined as in Eq. 23. If 1 and 1, 2 and 4 hold, then with probability at least $1-\delta$, it (jointly) holds that
@@
Proof of Lemma 10.
Recall from Algorithm 1 that we have defined
@@
Using the inequality $S_{t}\geq(t+1)t\sigma_{t}$, which holds thanks to 1, and applying Lemma 8 and Proposition 2, with probability at least $1-\delta$, we have
@@
Consider the sequence $\sigma_{t}:=\varsigma(t+1)^{-p}$ for $t\geq 0$ and the quantities defined in Eq. 23. If $p\in(0,1)$, then the sequence $\{\sigma_{t}\}_{t\geq 0}$ satisfies 1, 2 and 4 with $L=p$. Furthermore,
@@
$\quad+\frac{2}{1-p}\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\left(\frac{1}{2}+16\sqrt{\frac{2}{e(1-p)}+\log\left(\frac{16}{\delta}\right)}\right)$
$\quad+8\sqrt{2}\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\sqrt{\frac{1}{e(1-p)}+\log\left(\frac{16}{\delta}\right)}.$
Proof of Lemma 11.
We observe that $C_{q},V_{q}$ can be bounded from above by $C,V$ defined in (15) under the conditions in Lemma 6. Therefore, we focus on $\bar{C}_{\delta,q}$ for the remainder of the proof. Using the basic inequality $\log(x)/x\leq 1/e$ for any $x>0$, observe that for any $t\geq q$

### v2

@@ B.1 Proof of Theorem˜1
Proof of Theorem˜1.
The first claim on the non-asymptotic convergence guarantee follows directly from Lemmas˜7 and 8. As for the second claim on the asymptotic convergence, we prove a more general result: even if $\sigma_{t}$ is not set as $\varsigma(t+1)^{-p}$, the iterates of Algorithm˜1 still converge almost surely, provided the regularization sequence satisfies the conditions in Lemma˜6.
Lemma˜6 implies that $\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}$ with probability at least $1-\delta$. As $\delta$ can be arbitrarily small, we may conclude that $\lim_{t\to\infty}G(x_{t})=G_{\operatorname{opt}}$, almost surely. This implies that any limit point of $\{x_{t}\}_{t\geq 0}$ is in $X_{\operatorname{opt}}$ almost surely. Since $F$ is convex, hence lower semi-continuous over $X$, and by definition of $F_{\operatorname{opt}}$, we have $\liminf_{t\to\infty}F(x_{t})\geq F_{\operatorname{opt}}$ almost surely.
Besides, by combining Propositions˜1 and 2, we deduce that for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have, for all $t\geq 1$,
$F(x_{t})-F_{\operatorname{opt}}$ $\leq\frac{4c\left(4(\sigma_{0}L_{f}+L_{g})D^{2}+3(\sigma_{0}\sigma_{f}+\sigma_{g})D\right)}{(t+1)^{1/2}\sigma_{t}}\sqrt{\log\left({8dt^{2}}/{\delta}\right)}$
$\leq\frac{4c\left(4(\sigma_{0}L_{f}+L_{g})D^{2}+3(\sigma_{0}\sigma_{f}+\sigma_{g})D\right)}{(t+1)^{1/2}\sigma_{t}}\sqrt{\log\left({8dt^{2}}/{\delta}\right)}$
where the second inequality holds due to ˜1. We claim that all terms on the right hand side converge to 0 as $t\to\infty$. This is obvious except for the last term.
Recall that almost surely $\liminf_{t\to\infty}F(x_{t})\geq F_{\operatorname{opt}}$. Therefore we have $\limsup_{t\to\infty}(F_{\operatorname{opt}}-F(x_{t}))\leq 0$ and hence $\lim_{t\to\infty}\max\{F_{\operatorname{opt}}-F(x_{t}),0\}=0$ almost surely. Applying the Stolz–Cesàro theorem and (34) leads to
almost surely. Then each term on the right hand side of the inequality converges to 0 as $t\to\infty$. It follows that with probability at least $1-\delta$,
$\limsup_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})$ $\leq 0.$
Thus, since this holds for any $\delta\in(0,1)$, it holds almost surely. In conclusion, we have shown that $\limsup_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\leq 0$ and $\liminf_{t\to\infty}(F(x_{t})-F_{\operatorname{opt}})\geq 0$, which implies that $\lim_{t\to\infty}F(x_{t})=F_{\operatorname{opt}}$ almost surely. This completes the proof. ∎
B.2 Proof of Theorem˜2
We restate Theorem˜2 with explicit upper bounds on the stationary gaps and for a more general choice of $p$ and $\omega$. Toward the end of the proof, we show that the optimal parameters are $p=2/7$ and $\omega=6/7$.
Theorem 2.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜1 with stepsizes $\alpha_{t}=(t+1)^{-\omega}$ and regularization parameters $\sigma_{t}=\varsigma(t+1)^{-p}$ for any $t\geq 0$. If Assumption 1 holds with the exception that $F$ may be nonconvex, then with probability at least $1-\delta$, for every $t\geq 0$, it (jointly) holds that
$\frac{\sum_{i=0}^{t}(i+1)^{-p}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}(i+1)^{-p}}$
$\leq 4(1-p)\sup_{z\in X}|F(z)|(t+1)^{\omega-1}+2(1-p)F(x_{0})(t+1)^{p-1}+\tfrac{2(1-p)(\varsigma L_{f}+L_{g})D^{2}}{\varsigma(1-\omega)}(t+1)^{p-\omega}$
$+\tfrac{4c(1-p)\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{\varsigma(1-\omega/2)}(t+1)^{p-\omega/2}\sqrt{\log\left({4d(t+2)^{2}}/{\delta}\right)},$
and
$\quad+\frac{2c\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{1-\omega/2}t^{1-3\omega/2}\sqrt{\log\left({4d(t+q)^{2}}/{\delta}\right)},$
where $c$ is some absolute constant.
Proof of Theorem˜2.
$\quad+\frac{(\varsigma L_{f}+L_{g})D^{2}}{1-\omega}(t+1)^{1-\omega}.$
Applying Lemma˜3 with $k=0$, with probability at least $1-\delta$, we thus have
$\begin{aligned} \mathcal{G}(x_{t+1})&\!\leq\!2\varsigma\sup_{z\in X}|F(z)|(t\!+\!1)^{-p}\!+\!\varsigma F(x_{0})\!-\!\sum_{i=0}^{t}\sigma_{i}\mathcal{F}(x_{i})(t\!+\!1)^{-\omega}\!+\!\tfrac{(\varsigma L_{f}\!+\!L_{g})D^{2}}{1-\omega}(t\!+\!1)^{1\!-\!2\omega}\\ &\quad\!+\!\tfrac{2\sqrt{2}c(2(\varsigma L_{f}\!+\!L_{g})D\!+\!\frac{3^{\omega}}{3^{\omega}\!-\!1}(\varsigma\sigma_{f}\!+\!\sigma_{g}))D}{1\!-\!\omega/2}(t\!+\!1)^{1\!-\!3\omega/2}\sqrt{\log\left(\tfrac{4d(t\!+\!2)^{2}}{\delta}\right)}\end{aligned}$ (17)
Note that the maximum in the definition of function $\mathcal{F}$ is with respect to $X_{\operatorname{opt}}$, not $X$. Therefore, $\mathcal{F}(x_{i})$ can take both positive and negative values since $x_{i}\in X$. However, we can derive the following bound
which further implies that
$\quad+\frac{2c\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{1-\omega/2}t^{1-3\omega/2}\sqrt{\log\left(\frac{4d(t+q)^{2}}{\delta}\right)}.$ (18)
We next focus on the upper-level problem. Since $\mathcal{G}(x_{t+1})\geq 0$, using Eq.˜17, we have
$\quad+\frac{2c\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{1-\omega/2}(t+1)^{1-\omega/2}\sqrt{\log\left(\frac{4d(t+2)^{2}}{\delta}\right)}.$
we arrive at
$\frac{\sum_{i=0}^{t}\sigma_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\sigma_{i}}$ $\leq 4(1-p)\sup_{z\in X}|F(z)|(t+1)^{\omega-1}+2(1-p)F(x_{0})(t+1)^{p-1}$ (19)
$\quad+\tfrac{4c(1-p)\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{\varsigma(1-\omega/2)}(t+1)^{p-\omega/2}\sqrt{\log\left(\frac{4d(t+2)^{2}}{\delta}\right)}.$
Both bounds involving $\mathcal{G}$ and $\mathcal{F}$ hold with probability $\geq 1-\delta$. Hence, the first claim on the non-asymptotic convergence guarantee follows. Moreover, by optimizing over the parameters $p$ and $\omega$, we aim to ensure that the right-hand sides of both bounds converge to 0. To achieve this, we select $p$ and $\omega$ to minimize the slowest rate with respect to $t$:
which is realized by setting $p=2/7$, $\omega=6/7$ as required. Substituting these values yields the bound presented in the main body of the paper.
As for the second claim on asymptotic convergence, we prove a more general result: If
$\max\left\{-p,1-p-\omega,-\omega,1-2\omega,1-3\omega/2,\omega-1,p-1,p-\omega,p-\omega/2\right\}<0,$
then it holds that
$\liminf_{t\to\infty}\mathcal{F}(x_{t})=\lim_{t\to\infty}\mathcal{G}(x_{t})=0.$
Under this additional assumption and the fact that $\mathcal{G}(x_{t})\geq 0$, we deduce that $\lim_{t\to\infty}\mathcal{G}(x_{t})=0$ from taking the limit on both sides of (18). From the continuity of $G$, we deduce that any limit point of $\{x_{t}\}_{t\geq 0}$ is in $X_{\operatorname{opt}}$. the continuity of $\mathcal{F}$ from Lemma˜4 and the fact that $\mathcal{F}(x)\geq 0$ for any $x\in X_{\operatorname{opt}}$, we deduce that $\liminf_{t\to\infty}\mathcal{F}(x_{t})\geq 0$. Recall that
$\liminf_{t\to\infty}\mathcal{F}(x_{t})=\lim_{t\to\infty}\inf_{k\geq t}\mathcal{F}(x_{k}),$
and
$\frac{\sum_{i=0}^{t}\sigma_{i}\mathcal{F}(x_{i})}{\sum_{i=0}^{t}\sigma_{i}}\geq\frac{\sum_{i=0}^{t}\sigma_{i}\inf_{k\geq i}\mathcal{F}(x_{k})}{\sum_{i=0}^{t}\sigma_{i}}.$
Combining these observations and (19), we have
$\frac{\sum_{i=0}^{t}\sigma_{i}\inf_{k\geq i}\mathcal{F}(x_{k})}{\sum_{i=0}^{t}\sigma_{i}}$
$\leq 4(1-p)\sup_{z\in X}|F(z)|(t+1)^{\omega-1}+2(1-p)F(x_{0})(t+1)^{p-1}$
$\quad+\tfrac{4c(1-p)\sqrt{2}(2(\varsigma L_{f}+L_{g})D+\frac{3^{\omega}}{3^{\omega}-1}(\varsigma\sigma_{f}+\sigma_{g}))D}{\varsigma(1-\omega/2)}(t+1)^{p-\omega/2}\sqrt{\log\left(\frac{4d(t+2)^{2}}{\delta}\right)}.$
By the Stolz–Cesàro theorem (since $p\in(0,1)$, $\sum_{t\geq 0}\sigma_{t}=\infty$) and taking the limit on both sides of the above inequality, we obtain $\liminf_{t\to\infty}\mathcal{F}(x_{t})\leq 0$. Since this holds with probability at least $1-\delta$ for any $\delta\in(0,1)$, it should hold almost surely. Thus, we conclude the proof. ∎
Appendix C Proofs of Section 3
C.1 Proof of Theorem˜3
We restate Theorem˜3 with exact upper bounds on sub-optimality gaps.
Theorem  3.
Suppose the expectations defining $F$ and $G$ in Eq.˜1 are from uniform distributions over finite sets of size $[n]$. Let $\{z_{t}\}_{t\geq 1}$ be the iterates generated by Algorithm˜2 with $\alpha_{t}=\log(q)/q$ for $0\leq t<q$, $\alpha_{t}=2/(t+2)$ for any $t\geq q$, and $\sigma_{t}:=\varsigma(\max\{t,q\}+1)^{-p}$ for some chosen $\varsigma>0$, $p\in(0,1)$ and $S=q$. Under ˜1, with probability at least $1-\delta$, for any $t\geq q$, it holds that
$F(z_{t})-F_{\text{opt}}$
$\leq\frac{\max\left\{F(x_{0})-F_{\text{opt}},0\right\}+(G(x_{0})-G_{\text{opt}})/\varsigma+(L_{f}+L_{g}/\varsigma)D^{2}\left(\frac{1}{2}+16\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}\right)\log(q)}{(q+1)^{-1}(t+1)^{1-p}t}$
$\quad+\frac{2(L_{f}+L_{g}/\varsigma)D^{2}(t-q)+\left(8\sqrt{2}t-4q\right)(L_{f}+L_{g}/\varsigma)D^{2}\sqrt{\log\left(\frac{16t^{2}}{\delta}\right)}}{(t+1)^{1-p}t},$
$G(z_{t})-G_{\text{opt}}$
$\leq\left((1+2p)\left(F_{\text{opt}}-\min_{x\in X}F(x)\right)\varsigma+2(\varsigma L_{f}+L_{g})D^{2}+\left(\max_{x\in X}F(x)-f(x_{0})\right)\varsigma\right$.
$\quad\left.+\max_{x\in X}g(x)-g(x_{0})+\frac{2}{1-p}\left(L_{f}\varsigma+L_{g}\right)D^{2}\left(\frac{1}{2}+16\sqrt{\frac{2}{e(1-p)}+\log\left(16/\delta\right)}\right)\right$.
$\quad\left.+8\sqrt{2}\left(L_{f}\varsigma+L_{g}\right)D^{2}\sqrt{\frac{1}{e(1-p)}+\log\left(16/\delta\right)}\right)\left(1+\frac{2p}{\min\{1,2(1-p)\}}\right)(t+1)^{-p}.$
Moreover, with probability $1$, it holds that
@@ C.1 Proof of Theorem˜3
The proof proceeds by applying Proposition˜1 with $k=q$. Unlike the proof of Theorem˜1, here we have two main random components in Eq.˜10 and Eq.˜11: $(q+1)q\left(\Phi_{q}(x_{q})-\Phi_{q}(x_{\operatorname{opt}})\right)$ and $\sum_{i\in[t]\setminus[q]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$. We first derive probabilistic upper bounds for these terms, then carefully set the parameters in Algorithm˜2 to establish the convergence rate for both lower- and upper-level problems. In the following we use the notation
$F(x)=\frac{1}{n}\sum_{i\in[n]}f_{i}(x),\quad G(x)=\frac{1}{n}\sum_{i\in[n]}g_{i}(x).$
@@
We begin with a probabilistic bound on the gradient estimator used in Algorithm˜2.
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with $\alpha_{t}=\log(q)/q$ for any $0\leq t<q$ and $\alpha_{t}=2/(t+2)$ for any $t\geq q$. If ˜1 holds, then for any $t\geq 0$, given $\delta\in(0,1)$, with probability at least $1-\delta$, it (jointly) holds that
$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|$ $\leq 8L_{f}D\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(8/\delta\right)},$ (20)
$\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|$ $\leq 8L_{g}D\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(8/\delta\right)},$
@@
$\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|$ $\leq 8L_{f}D\sqrt{\frac{(t-s_{t})}{S(t+1)(s_{t}+1)}}\sqrt{\log\left(8/\delta\right)}$ (21)
$\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|$ $\leq 8L_{g}D\sqrt{\frac{(t-s_{t})}{S(t+1)(s_{t}+1)}}\sqrt{\log\left(8/\delta\right)},$
@@
if $t>q$. Given $\delta\in(0,1)$ and $1\leq t\leq q$, setting the right hand side $\delta/2$ and solving for $\lambda$, we have
$P\left(\left\|\widehat{\nabla F}_{t}-\nabla F(x_{t})\right\|\geq 8L_{f}D\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(8/\delta\right)}\right)\leq\frac{\delta}{2}$
@@
$P\left(\left\|\widehat{\nabla G}_{t}-\nabla G(x_{t})\right\|\geq 8L_{g}D\frac{\log(q)}{\sqrt{qS}}\sqrt{\log\left(8/\delta\right)}\right)\leq\frac{\delta}{2}.$
@@
To establish convergence results for Algorithm˜2, we replace ˜3 with the following.
@@
Note that $\sigma_{t}$ is fixed during the initialization phase, and $\sigma_{t}^{2}$ in ˜3 is replaced by $\sigma_{t}^{2}$ in ˜4. Unlike the proof of Theorem˜1, where we set $k=0$, the proof of Theorem˜3 requires setting $k=q$. We therefore begin by analyzing the sequence $\{x_{t}\}_{t\geq 0}$ generated by Algorithm˜2 over the first $q$ iterations.
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with $\alpha_{t}=\log(q)/q$ for any $0\leq t<q$ and $S=q$. If ˜1 and ˜4 hold, we have then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{0\leq i<q}1/(i+1)(i+2)$, for any $t\geq q$, it holds that
@@
Proof of Lemma˜10.
By Lemma˜1, for any $0\leq i<q$, we have
$\begin{aligned} \Phi_{i}(x_{i+1})-\Phi_{i}(x_{\operatorname{opt}})&\leq(1-\alpha_{i})(\Phi_{i}(x_{i})-\Phi_{i}(x_{\operatorname{opt}}))+\frac{(L_{f}\sigma_{i}+L_{g})D^{2}\alpha_{i}^{2}}{2}\\ &\quad+2D\alpha_{i}\|\nabla\Phi_{i}(x_{i})-\widehat{\nabla\Phi}_{i}\|.\end{aligned}$ (22)
By ˜4, we have $\alpha_{0}=\dots=\alpha_{q-1}$ and
@@
where both inequalities are direct consequences of ˜4 and (22). Note that
@@
$\begin{aligned} \Phi_{q}(x_{q})-\Phi_{q}(x_{\operatorname{opt}})&\leq(1-\alpha_{0})^{q}\left(\Phi_{0}(x_{0})-\Phi_{0}(x_{\operatorname{opt}})\right)+\sum_{i=0}^{q-1}\frac{(L_{f}\sigma_{0}+L_{g})D^{2}\alpha_{0}}{2}\\ &\quad+\sum_{i=0}^{q-1}2D\|\nabla\Phi_{i}(x_{i})-\widehat{\nabla\Phi}_{i}\|.\end{aligned}$ (23)
@@
Hence, applying Lemma˜9 along with a union bound argument implies that, for any $\delta\in(0,1)$, with probability at least $1-\delta\sum_{0\leq i<q}1/(i+1)(i+2)$, we have
@@
Plugging the above probabilistic bound into Eq.˜23, noting that $t\geq q$ and $S=q$, and using the definition of $\alpha_{0}$ complete the proof. ∎
We next analyze the sequence $\{x_{t}\}_{t\geq 0}$ generated by Algorithm˜2 after the first $q$ iterations.
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with $\alpha_{t}=\log(q)/q$ for any $0\leq t<q$, $\alpha_{t}=2/(t+2)$ for any $t\geq q$ and $S=q$. If ˜1 holds and the parameters $\{\sigma_{t}\}_{t\geq 0}$ are non-increasing and positive, then given $\delta\in(0,1)$, with probability at least $1-\delta\sum_{i\in[t]\setminus[q]}1/i(i+1)$, for any $t>q$, it holds that
@@
Given $\delta\in(0,1)$ and $q<i\leq 2q$, by Lemma˜9 and the inequality Eq.˜25, with probability at least $1-\delta/i(i+1)$, we have
@@
where the second inequality holds as $\frac{1}{q}\sum_{q<i\leq t}i\leq\sqrt{2}t-q/2$ for all $q<t\leq 2q$. Similarly, given $\delta\in(0,1)$ and $i>2q$, by Lemma˜9 and the inequality Eq.˜25, with probability at least $1-\delta/i(i+1)$, we have
@@
where the second inequality follows since $q\lfloor(i-1)/q\rfloor+1\geq i-q$, and the third inequality follows since i/(iq)<2i/(i-q)<2 when $i>2q$. Thus, if $t>2q$, with probability at least $1-\delta\sum_{2q<i\leq t}1/i(i+1)=1-\delta$, we have
@@
Lemma˜17 in Section˜D.1 guarantees that these quantities remain finite for any parameter choice satisfying 1 to 2 and 4. In particular, these conditions hold when $\sigma_{t}=\varsigma(t+1)^{p}$ for any $p\in(0,1)$, which is what is used in Theorem˜3. We now analyze the sequence $\{x_{t}\}_{t\geq 0}$ in Algorithm˜2 for the inner-level problem.
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with $\alpha_{t}=\log(q)/q$ for $0\leq t<q$ $\alpha_{t}=2/(t+2)$ for any $t\geq q$, $S=q$, and let $C_{\delta,q}$ be defined as in (26g). If ˜1 and ˜1, 2 and 4 hold, then with probability at least $1-\delta$, it holds that
@@
Proof of Lemma˜11.
Combining (10) from Proposition˜1 with the probabilistic bounds in Lemma˜10 and Proposition˜3, and after some straightforward calculations, with probability at least $1-\delta$, we have
@@
We next analyze the convergence of the sequence $\{z_{t}\}_{t\geq 0}$ in Algorithm˜2 for both outer- and inner-level problems.
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜2 with $\alpha_{t}=\log(q)/q$ for $0\leq t<q$ $\alpha_{t}=2/(t+2)$ for any $t\geq q$, $S=q$, and let $C_{\delta,q}$ and $V_{q}$ defined as in Eq.˜26. If ˜1 and ˜1, 2 and 4 hold, then with probability at least $1-\delta$, it (jointly) holds that
@@
Proof of Lemma˜12.
Recall from Algorithm˜1 that we have defined
@@
Using the inequality $S_{t}\geq(t+1)t\sigma_{t}$, which holds thanks to ˜1, and applying Lemma˜10 and Proposition˜2, with probability at least $1-\delta$, we have
@@
Consider the sequence $\sigma_{t}:=\varsigma(t+1)^{-p}$ for $t\geq 0$ and the quantities defined in Eq.˜26. If $p\in(0,1)$, then the sequence $\{\sigma_{t}\}_{t\geq 0}$ satisfies ˜1, 2 and 4 with $L=p$. Furthermore,
@@
$\quad+\frac{2}{1-p}\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\left(\frac{1}{2}+16\sqrt{\frac{2}{e(1-p)}+\log\left(16/\delta\right)}\right)$
$\quad+8\sqrt{2}\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\sqrt{\frac{1}{e(1-p)}+\log\left(16/\delta\right)}.$
Proof of Lemma˜13.
We observe that $C_{q},V_{q}$ can be bounded from above by $C,V$ defined in (15) under the conditions in Lemma˜8. Therefore, we focus on $\bar{C}_{\delta,q}$ for the remainder of the proof. Using the basic inequality $\log(x)/x\leq 1/e$ for any $x>0$, observe that for any $t\geq q$